how much of 75% copper alloy should be melted into 62 kg of a 35% copper alloy to produce an alloy which is 50 % copper?
Answer by mananth(16946)

-------------- percent ---------------- Amount
Alloy I 35 ---------------- 62 kg
Alloy II 75 ---------------- x kg
Mixture 50 ---------------- 62 + x kg

35 * 62 + 75 x = 50 ( 62 + x )
2170 + 75 x = 3100 + 50 x
75 x -50 x = -2170 + 3100
25 x = 930
/ 25
x = 37.2 kg Alloy II

m.ananth@hotmail.ca